# Elasticity without a reference state: continuum mechanics of active tension nets

### v1

@@ Abstract
A constitutive relation between stress and strain relative to a reference state is the basic assumption of elasticity theory. However, in living matter, stress is governed by (motor molecule) activity rather than a constitutive law. What paradigm takes the place of elasticity in this setting? Here, we derive a continuum theory of active mechanics by taking the continuum limit of the Active Tension Network model of 2d epithelia. Instead of a reference state, we start from a prescribed active force configuration, encoded in a Riemannian “tension metric”. Intuitively, one expects cells to adjust their positions to achieve force balance by rearranging local sources of active stress. More precisely, the cell positions define an embedding of the tension metric into 2d physical space, which determines the macroscopic physical stress. For free boundaries, tissue adopts a certain intrinsically defined shape, the force-balanced embedding with minimal internal stress. Boundary forces then deform this embedding. The resulting stress transformation yields an effective stress-strain relation. Key elements of elasticity hence emerge from a “stress-only” starting point, explaining how tissue shape can be adiabatically controlled by active stress during morphogenesis. Plastic behavior arises from topological cell rearrangement, which we represent by a continuous reparameterization of the tension metric, providing a principled continuum theory of emergent elasto-plastic flow. To express this physics, we use the mathematics of isothermal coordinates and quasi-conformal maps. The present theory elucidates the unconventional mechanics of living tissues and may apply to 2d active and granular materials more generally.
@@ Introduction
During morphogenesis, cells sculpt the physical form of the developing embryo. In contrast to inert materials, living tissues generate forces internally [1, 2] Key challenges are to characterize the macroscopic mechanical properties (i.e. the “phase of matter”) of developing tissues and to understand the control of tissue shape (morphogenesis) by the cell-generated forces.
@@
Here, we derive a continuum theory of emergent, active elasticity. The key to our approach lies in representing the configuration of active stress by a Riemannian metric $\mathbf{g}$, a continuum version of Maxwell–Cremona force tesselations [14]. The cell tiling defines an embedding of the tension metric into physical space. We show that the central elements of elasticity theory, a stress-free reference state and a stress-strain relationship, emerge from this embedding. Additionally, we will demonstrate that cell-scale topological rearrangements that enable large deformations can be captured in the continuum theory via a suitable reparametrization of the tension metric. Taken together, these ideas enable a continuum description of adiabatic tissue flow driven by active stress dynamics. The resulting theory is a form of “emergent” viscoelasticity, in which adiabatic stress dynamics can drive active plastic flow in the absence, or against the resistance, of external forces. In a forthcoming companion [15], we will show that the continuum-based, “top-down” approach presented here precisely matches a “bottom-up” coarse-graining analysis of the ATN model.
@@ Results
Maps satisfying the Cauchy–Riemann equation $\bar{\partial}_{z}f=0$, i.e. $z\mapsto z^{\prime}=f(z)$, are conformal (angle-preserving). Their Jacobian is a diagonal matrix ${\mathbf{D}\kern-0.80002ptf}=\mathrm{diag}[\partial_{z}f,\overline{\partial_{z}f}]$, combining rotation and scaling by the conformal factor $\lambda_{z|z^{\prime}}=|\partial_{z}f|$. Furthermore, in this case, $\Delta\log\lambda_{z|z^{\prime}}=4\partial_{z}\bar{\partial}_{z}\operatorname{Re}[\log(\partial_{z}f)]=0$. More generally, the shear created by a map is measured by its Beltrami coefficient $\mu_{z|z^{\prime}}=\bar{\partial}_{z}f/\partial_{z}f$ (the subscript of $\mu_{z|z^{\prime}}$ denotes the map’s source and target.) Note that the Beltrami coefficients of the map $f$ and the quadratic form $({\mathbf{D}\kern-0.80002ptf})^{\dagger}\cdot{\mathbf{D}\kern-0.80002ptf}$ agree; $()^{\dagger}$ being the Hermitian adjoint. Geometrically, $f$ maps infinitesimal circles to ellipses with eccentricity $|\mu_{z|z^{\prime}}|/(1-|\mu_{z|z^{\prime}}|^{2})^{1/2}$. If $|\mu_{z|z^{\prime}}|<1$ everywhere, $f$ is called quasi-conformal (QC). In particular, QC maps preserve orientation. Notation-wise, their argument distinguishes QC maps $f(z,\bar{z})$ (also written $f(\mathbf{z})$), from conformal maps $f(z)$. In the following, we use the formalism of QC maps to develop our continuum theory.
@@
${}K_{g}=-\lambda_{g}^{-2}\big($ $\Delta\log\lambda_{g}+2\operatorname{Re}[\partial_{\zeta}\bar{\partial}_{\zeta}dg_{\bar{z}z}+\partial_{z}^{2}dg_{zz}]\big)$ (13)
@@
An analogous, if lengthy, calculation shows that for $K_{g}\neq 0$, $z=\lambda_{g}^{-2}\bar{\partial}_{w}\theta$ is a perturbative solution to Eq. (24) (it corresponds to the Riemannian gradient of $\theta$ with respect to $\mathbf{g}$). The manifold of force-balanced states parametrized by $\theta$ corresponds to the “isogonal” modes of the discrete ATN model [21, 11], and we will therefore refer to $\theta$ as the isogonal potential (this connection will be detailed in the companion paper [15]). Conformal and isogonal displacements exhaust all solutions to Eq. (24) (Appendix B). Any physical configuration $w$ can hence be obtained from the reference $z$ by a composition of a potential map and a conformal map. The two modes suffice to accommodate arbitrary deformations imposed on the system boundary (Appendix C).
@@
To understand the physics resulting from the interplay of conformal–isogonal structure and the constitutive law for pressure, we now expand stress and pressure for small deformations $w=z+u(z,\bar{z}),\;|\partial_{z}u|\ll 1$ around the isothermal, stress-free reference state. We further assume small cell-density gradients in the reference state, $n(z,\bar{z})=n_{0}(1+\delta n(z,\bar{z}))$ (this requires that the tension curvature $K_{g}$, and hence $\log\lambda_{g}$, are also small). In the discrete setting, cell areas are close to the Voronoi areas of the tension triangulation.
@@
A remarkable consequence of these results is that the shape and macroscopic mechanical properties of the tissue are determined by the tension metric and cell density alone. There are many microscopic tension configurations (tension triangulations) that have the same $\mathbf{g}$ and $n$. On a macroscopic level, they are all equivalent – the microscopic details are irrelevant for the static problem. Moreover, up to the overall scale $p_{0}$, the only free parameter of the theory is the cellular bulk modulus $B$. Let us briefly discuss four special cases: (i) For incompressible cells, $B\rightarrow\infty$ forcing $\partial_{\alpha}u_{\alpha}=-\delta n$. (ii) If pressures follow the ideal gas law $p(n)p_{0}n/n_{0}$, i.e. $B=1$, the total bulk and shear moduli are equal, which may be relevant to experiments on conventional foams with air-filled bubbles. (iii) For $B=1/2$, the total bulk modulus vanishes, so that the conformal mode $f_{\alpha}$ becomes soft. If $B<1/2$, the total bulk modulus is negative: intracellular pressure is insufficient to balance the isotropic tensile stress. Mechanical stability then requires a constraint on the total area. (iv) For fully compressible cells $B=0$, the isogonal potential $\theta$ is an unconstrained soft mode. However, conformal modes, which require pressure gradients, cannot be accommodated. By Eq. (36), freely compressible cells (i.e., $B=0$) are unstable if $K_{g}\neq 0$.
@@
$=\frac{\mu_{\zeta|z}+e^{-i\phi_{\zeta|z}}\mu_{z|w}}{1+e^{-i\phi_{\zeta|z}}\bar{\mu}_{\zeta|z}\mu_{z|w}}\approx\mu_{\zeta|z}+e^{-i\phi_{\zeta|z}}\mu_{z|w}$ (48)
where $e^{-i\phi_{\zeta|z}}=\overline{\partial_{\zeta}z}/\partial_{\zeta}z$, and we assumed $|\mu_{\zeta|z}|,|\mu_{z|w}|\ll 1$ to make the approximation in the last step.
@@
$\approx\gamma_{\mathrm{T1}}(\mu_{\zeta|z}+e^{-i\phi_{\zeta|z}}\mu_{z|w})$ (55)
@@
$\left(\partial_{t}\mu_{g}(\zeta,\bar{\zeta})\right)_{\mathrm{T1}}=\gamma_{\mathrm{T1}}e^{-i\phi_{z}}\mu_{z|w}=-\gamma_{\mathrm{T1}}\frac{\delta|\mu_{z|w}|^{2}}{\delta\bar{\mu}_{g}},$ (58)
@@ Discussion
To describe tissue flow driven by adiabatic dynamics of active stress, one tracks the reference state $z(\xi,\bar{\xi},t)$ of the evolving tension metric $\mathbf{g}(t)$. The time evolution $\dot{z}$ is governed by a Beltrami flow, Eq. (40). Large deformations entail topological cell rearrangement. This raises the challenge of incorporating discrete topology into a continuum theory. On a cell level, topological information can be “geometrized” via the so-called circle packing of the cell adjacency graph, whose existence and uniqueness is guaranteed by the Koebe–Andreev–Thurston theorem [33]. As we will detail in the companion paper [15], the circle packing is a “discrete” conformal embedding of the equilateral triangulation formed by the adjacency graph. In the continuum, it becomes the isothermal embedding $\zeta(\xi,\bar{\xi})$ of the adjacency metric $\mathbf{a}$. Tension dynamics, plastic flow, and elastic deformation can thus be formulated as two coupled Beltrami flows, Eqs. (40), and (52). This principled approach naturally incorporates the local tension configuration via the Jacobian $\partial_{\bm{\zeta}}\mathbf{z}$.
@@
The link between the continuum theory and cell level (the discrete ATN model) will be discussed in detail in a companion paper [15]. There, we exploit a theory of discrete conformal maps [45, 46] to bridge the gap between the continuum theory of the present manuscript and the mechanics at the cellular level. This theory constructs a discrete equivalent of the isothermal embedding $z(\xi,\bar{\xi})$. The isogonal (curl-free) and conformal modes identified by the continuum analysis reappear as the geometric soft modes which respect the microscopic force balance constraints [47, 11, 48], while the Liouville equation for the curvature of the tension manifold leads to a generalized von Neumann law for the cell pressures. Coarse-graining the microscopic stress confirms the stress-transformation law Eq. (21). Thus, the “top-down” analysis presented here is independently confirmed by “bottom-up” coarse-graining.
@@
While we have introduced our theory in the context of epithelial tissue mechanics, we believe it applies more broadly. Non-epithelial tissues, like those composed of protrusive, mesenchymal cells, do not form tension nets in the strict ATN sense, but their architecture and shape are also controlled by balancing active contractile stresses against bulk elasticity. Beyond the biological setting, 2D fluid or ferro-fluid foams [49, 13] correspond to the limiting case where all $\tau_{ij}=\tau=\mathrm{const}$. Tension dynamics could be introduced, for instance, through chemical reactions within foam bubbles. As explained in a forthcoming companion paper [15], there is also a close connection to granular materials and truss networks, which can be understood as the mechanical Legendre dual of tension nets. In granular materials, one typically studies the ensemble of compatible (force-balanced) stress configurations for a given physical arrangement of particles. This is, in a sense, the dual of the problem considered in the present study. In future work, the Legendre duality might, therefore, provide a starting point for building a continuum theory for granular matter. Indeed, previous phenomenological approaches postulate that granular materials can also be described by “emergent elasticity” [50].
@@ Appendix F Kinematics of Beltrami flow
$\mu_{\zeta^{\prime}|z}$ $(\zeta^{\prime},\bar{\zeta}^{\prime})=\frac{\mu_{\zeta^{\prime}|\zeta}+\mu_{\zeta|z}\,e^{-i\phi_{\zeta^{\prime}|\zeta}}}{1+\bar{\mu}_{\zeta^{\prime}|\zeta}\,\mu_{\zeta|z}\,e^{-i\phi_{\zeta^{\prime}|\zeta}}}$ (75)
@@
where the phase factor is $e^{-i\phi_{\zeta^{\prime}|\zeta}}=(1-\overline{\partial_{\zeta}{\dot{\zeta}}})/(1-\partial_{\zeta}{\dot{\zeta}})\approx 1+i2\mathrm{Im}[\partial_{\zeta}{\dot{\zeta}}]dt$. The displacement change from $\zeta$ to $\zeta^{\prime}$ leads to an additional advective term $\mathrm{Re}[\bar{\nu}\partial_{\zeta}]\mu_{z}dt$. Taken together, one arrives at

### v2

@@ Abstract
A constitutive relation between stress and strain relative to a reference state is the basic assumption of elasticity theory. However, in living matter, stress is governed by (motor molecule) activity rather than a constitutive law. What paradigm takes the place of elasticity in this setting? Here, we derive a continuum theory of active mechanics by taking the continuum limit of the Active Tension Network model of 2d epithelia. Instead of a reference state, we start from a prescribed active force configuration, encoded in a Riemannian “tension metric”. Intuitively, one expects cells to adjust their positions to achieve force balance by rearranging local sources of active stress. More precisely, the cell positions define an embedding of the tension metric into 2d physical space, which determines the macroscopic physical stress. For free boundaries, tissue adopts a certain intrinsically defined shape, the force-balanced embedding with minimal internal stress. Boundary forces then deform this embedding. The resulting stress transformation yields an effective stress-strain relation. Key elements of elasticity hence emerge from a “stress-only” starting point, explaining how tissue shape can be adiabatically controlled by active stress during morphogenesis. Plastic behavior arises from topological cell rearrangement, which we represent by a continuous reparameterization of the tension metric, providing a principled continuum theory of emergent elasto-plastic flow. To express this physics, we use the mathematics of isothermal coordinates and quasi-conformal maps. An explicit coarse-graining study, described in a companion paper, substantiates the results of our continuum analysis. The present theory elucidates the unconventional mechanics of living tissues and may apply to 2d active and granular materials more generally.
@@ Introduction
During morphogenesis, cells sculpt the physical form of the developing embryo. In contrast to inert materials, living tissues generate forces internally [1, 2]. Key challenges are to characterize the macroscopic mechanical properties (i.e. the “phase of matter”) of developing tissues and to understand the control of tissue shape (morphogenesis) by the cell-generated forces.
@@
Here, we derive a continuum theory of emergent, active elasticity. The key to our approach lies in representing the configuration of active stress by a Riemannian metric $\mathbf{g}$, a continuum version of Maxwell–Cremona force tesselations [14]. The cell tiling defines an embedding of the tension metric into physical space. We show that the central elements of elasticity theory, a stress-free reference state and a stress-strain relationship, emerge from this embedding. Additionally, we will demonstrate that cell-scale topological rearrangements that enable large deformations can be captured in the continuum theory via a suitable reparametrization of the tension metric. Taken together, these ideas enable a continuum description of adiabatic tissue flow driven by active stress dynamics. The resulting theory is a form of “emergent” viscoelasticity, in which adiabatic stress dynamics can drive active plastic flow in the absence, or against the resistance, of external forces. In a companion paper [15], we show that the continuum-based, “top-down” approach presented here precisely matches a “bottom-up” coarse-graining analysis of the ATN model.
@@ Results
Maps satisfying the Cauchy–Riemann equation $\bar{\partial}_{z}f=0$, i.e. $z\mapsto z^{\prime}=f(z)$, are conformal (angle-preserving). Their Jacobian is a diagonal matrix ${\mathbf{D}\kern-0.80002ptf}=\mathrm{diag}[\partial_{z}f,\overline{\partial_{z}f}]$, combining rotation and scaling by the conformal factor $\lambda_{z|z^{\prime}}=|\partial_{z}f|$. Furthermore, in this case, $\Delta\log\lambda_{z|z^{\prime}}=4\partial_{z}\bar{\partial}_{z}\operatorname{Re}[\log(\partial_{z}f)]=0$. More generally, the shear created by a map is measured by its Beltrami coefficient $\mu_{z|z^{\prime}}=\bar{\partial}_{z}f/\partial_{z}f$ (the subscript of $\mu_{z|z^{\prime}}$ denotes the map’s source and target.) Note that the Beltrami coefficients of the map $f$ and the quadratic form $({\mathbf{D}\kern-0.80002ptf})^{\dagger}\cdot{\mathbf{D}\kern-0.80002ptf}$ agree; $()^{\dagger}$ being the Hermitian adjoint. Geometrically, $f$ maps infinitesimal circles to ellipses with eccentricity $|\mu_{z|z^{\prime}}|/(1-|\mu_{z|z^{\prime}}|^{2})^{1/2}$. If $|\mu_{z|z^{\prime}}|<1$ everywhere, $f$ is called quasi-conformal (QC). In particular, QC maps preserve orientation. Notation-wise, their argument distinguishes QC maps $f(z,\bar{z})$ (also written $f(\mathbf{z})$), from conformal maps $f(z)$. In the following, we use the formalism of QC maps to develop our continuum theory (in Ref. [18], a similar QC approach was used to describe the growth of 2d tissues).
@@
${}K_{g}=-\lambda_{g}^{-2}\big($ $\Delta\log\lambda_{g}+2\operatorname{Re}[\partial_{z}\bar{\partial}_{z}dg_{\bar{z}z}+\partial_{z}^{2}dg_{zz}]\big)$ (13)
@@
An analogous, if lengthy, calculation shows that for $K_{g}\neq 0$, $z=\lambda_{g}^{-2}\bar{\partial}_{w}\theta$ is a perturbative solution to Eq. (24) (it corresponds to the Riemannian gradient of $\theta$ with respect to $\mathbf{g}$). The manifold of force-balanced states parametrized by $\theta$ corresponds to the “isogonal” modes of the discrete ATN model [22, 11], and we will therefore refer to $\theta$ as the isogonal potential (this connection is detailed in the companion paper [15]). Conformal and isogonal displacements exhaust all solutions to Eq. (24) (Appendix B). Any physical configuration $w$ can hence be obtained from the reference $z$ by a composition of a potential map and a conformal map. The two modes suffice to accommodate arbitrary deformations imposed on the system boundary (Appendix C).
@@
To understand the physics resulting from the interplay of conformal–isogonal structure and the constitutive law for pressure, we now expand stress and pressure for small deformations $w=z+u(z,\bar{z}),\;|\partial_{z}u|\ll 1$ around the isothermal, stress-free reference state. We further assume small cell-density gradients in the reference state, $n(z,\bar{z})=n_{0}(1+\delta n(z,\bar{z}))$, and small curvature $K_{g}$ (hence, $\log\lambda_{g}$ is also small). In the discrete setting, cell areas are close to the Voronoi areas of the tension triangulation.
@@
A remarkable consequence of these results is that the shape and macroscopic mechanical properties of the tissue are determined by the tension metric and cell density alone. There are many microscopic tension configurations (tension triangulations) that have the same $\mathbf{g}$ and $n$. On a macroscopic level, they are all equivalent – the microscopic details are irrelevant for the static problem. Moreover, up to the overall scale $p_{0}$, the only free parameter of the theory is the cellular bulk modulus $B$.
Let us briefly discuss four special cases: (i) For incompressible cells, $B\rightarrow\infty$ forcing $\partial_{\alpha}u_{\alpha}=-\delta n$. (ii) If pressures follow the ideal gas law $p(n)p_{0}n/n_{0}$, i.e. $B=1$, the total bulk and shear moduli are equal, which may be relevant to experiments on conventional foams with air-filled bubbles. (iii) For $B=1/2$, the total bulk modulus vanishes, so that the conformal mode $f_{\alpha}$ becomes soft. If $B<1/2$, the total bulk modulus is negative: intracellular pressure is insufficient to balance the isotropic tensile stress. Mechanical stability then requires a constraint on the total area. (iv) For fully compressible cells $B=0$, the isogonal potential $\theta$ is an unconstrained soft mode. However, conformal modes, which require pressure gradients, cannot be accommodated. By Eq. (36), freely compressible cells (i.e., $B=0$) are unstable if $K_{g}\neq 0$.
@@
$=\frac{\mu_{\zeta|z}+e^{-i2\phi_{\zeta|z}}\mu_{z|w}}{1+e^{-i2\phi_{\zeta|z}}\bar{\mu}_{\zeta|z}\mu_{z|w}}\approx\mu_{\zeta|z}+e^{-i2\phi_{\zeta|z}}\mu_{z|w}$ (48)
where $e^{-i2\phi_{\zeta|z}}=\overline{\partial_{\zeta}z}/\partial_{\zeta}z$, and we assumed $|\mu_{\zeta|z}|,|\mu_{z|w}|\ll 1$ to make the approximation in the last step.
@@
$\approx\gamma_{\mathrm{T1}}(\mu_{\zeta|z}+e^{-i2\phi_{\zeta|z}}\mu_{z|w})$ (55)
@@
$\left(\partial_{t}\mu_{g}(\zeta,\bar{\zeta})\right)_{\mathrm{T1}}=\gamma_{\mathrm{T1}}e^{-i2\phi_{\zeta|z}}\mu_{z|w}=-\gamma_{\mathrm{T1}}\frac{\delta|\mu_{z|w}|^{2}}{\delta\bar{\mu}_{g}},$ (58)
@@ Discussion
To describe tissue flow driven by adiabatic dynamics of active stress, one tracks the reference state $z(\xi,\bar{\xi},t)$ of the evolving tension metric $\mathbf{g}(t)$. The time evolution $\dot{z}$ is governed by a Beltrami flow, Eq. (40). Large deformations entail topological cell rearrangement. This raises the challenge of incorporating discrete topology into a continuum theory. On a cell level, topological information can be “geometrized” via the so-called circle packing of the cell adjacency graph, whose existence and uniqueness is guaranteed by the Koebe–Andreev–Thurston theorem [34]. As explained in the companion paper [15], the circle packing is a “discrete” conformal embedding of the equilateral triangulation formed by the adjacency graph. In the continuum, it becomes the isothermal embedding $\zeta(\xi,\bar{\xi})$ of the adjacency metric $\mathbf{a}$. Tension dynamics, plastic flow, and elastic deformation can thus be formulated as two coupled Beltrami flows, Eqs. (40), and (52). This principled approach naturally incorporates the local tension configuration via the Jacobian $\partial_{\bm{\zeta}}\mathbf{z}$.
@@
The link between the continuum theory and cell level (the discrete ATN model) is discussed in detail in a companion paper [15]. There, we exploit a theory of discrete conformal maps [46, 47] to bridge the gap between the continuum theory of the present manuscript and the mechanics at the cellular level. This theory constructs a discrete equivalent of the isothermal embedding $z(\xi,\bar{\xi})$. The isogonal (curl-free) and conformal modes identified by the continuum analysis reappear as the geometric soft modes which respect the microscopic force balance constraints [48, 11, 49], while the Liouville equation for the curvature of the tension manifold leads to a generalized von Neumann law for the cell pressures. Coarse-graining the microscopic stress confirms the stress-transformation law Eq. (21). Thus, the “top-down” analysis presented here is independently confirmed by “bottom-up” coarse-graining.
@@
While we have introduced our theory in the context of epithelial tissue mechanics, we believe it applies more broadly. Non-epithelial tissues, like those composed of protrusive, mesenchymal cells, do not form tension nets in the strict ATN sense, but their architecture and shape are also controlled by balancing active contractile stresses against bulk elasticity. Beyond the biological setting, 2D fluid or ferro-fluid foams [50, 13] correspond to the limiting case where all $\tau_{ij}=\tau=\mathrm{const}$. Tension dynamics could be introduced, for instance, through chemical reactions within foam bubbles. As explained in the companion paper [15], there is also a close connection to granular materials and truss networks, which can be understood as the mechanical Legendre dual of tension nets. In granular materials, one typically studies the ensemble of compatible (force-balanced) stress configurations for a given physical arrangement of particles. This is, in a sense, the dual of the problem considered in the present study. In future work, the Legendre duality might, therefore, provide a starting point for building a continuum theory for granular matter. Indeed, previous phenomenological approaches postulate that granular materials can also be described by “emergent elasticity” [51].
@@ Appendix F Kinematics of Beltrami flow
$\mu_{\zeta^{\prime}|z}$ $(\zeta^{\prime},\bar{\zeta}^{\prime})=\frac{\mu_{\zeta^{\prime}|\zeta}+\mu_{\zeta|z}\,e^{-i2\phi_{\zeta^{\prime}|\zeta}}}{1+\bar{\mu}_{\zeta^{\prime}|\zeta}\,\mu_{\zeta|z}\,e^{-i2\phi_{\zeta^{\prime}|\zeta}}}$ (75)
@@
where the phase factor is $e^{-i2\phi_{\zeta^{\prime}|\zeta}}=(1-\overline{\partial_{\zeta}{\dot{\zeta}}})/(1-\partial_{\zeta}{\dot{\zeta}})\approx 1+i2\mathrm{Im}[\partial_{\zeta}{\dot{\zeta}}]dt$. The displacement change from $\zeta$ to $\zeta^{\prime}$ leads to an additional advective term $\mathrm{Re}[\bar{\nu}\partial_{\zeta}]\mu_{z}dt$. Taken together, one arrives at